OpenStudy (anonymous):

find the series solutions for y''-xy'+2y=0

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

this is a second order homogenous ODE \[ y''-xy'+2y=0 \] this probably turn out to be a series solution

OpenStudy (anonymous):

Yes but I'm not sure what those series are :(

OpenStudy (anonymous):

oh duh.. the question itself says about a series solution!!

OpenStudy (anonymous):

ok lets say the solution is: \[y(x)=\sum_{n=0}^\infty a_nx^n\\ \text{differentiating once,}\\ y'=\sum_{n=1}^\infty na_nx^{n-1}\\ \text{differentiating again,}\\ y''=\sum_{n=2}^\infty n(n-1)a_nx^{n-2} \]we now plug this in the given DE \[ \sum_{n=2}^\infty n(n-1)a_nx^{n-2}-x\times \sum_{n=1}^\infty na_nx^{n-1}+2\times\sum_{n=0}^\infty a_nx^n=0\\ \sum_{n=2}^\infty n(n-1)a_nx^{n-2}-\sum_{n=1}^\infty na_nx^n+\sum_{n=0}^\infty 2a_nx^n=0\\ \sum_{\color{red}{n=0}}^\infty (n+2)(n+1)a_{n}x^{n+1}-\sum_{\color{red}{n=0}}^\infty (n+1)a_{n+1}x^{n+1}+\sum_{n=0}^\infty 2a_nx^n=0\\ \sum_{n=0}^\infty\left[(n+2)(n+1)a_{n+2}x^{n}-(n+1)a_{n+1}x^{n+1}+2a_nx^n\right]=0\\ \implies\Large (n+2)(n+1)a_{n+2}-(n+1)a_{n+1}x+2a_n=0\\ \]

OpenStudy (anonymous):

yep i've gotten that far

OpenStudy (anonymous):

looks like I messed up there should not be any "x" in the recurrance relation \[ \sum_{\color{red}{n=1}}^\infty (n+1)(n)a_{n+1}x^{n+1}-\sum_{\color{red}{n=1}}^\infty na_{n}x^{n}+2a_0+\sum_{\color{red}{n=1}}^\infty 2a_nx^n=0\\ \sum_{\color{red}{n=1}}^\infty (n+1)(n)a_{n+1}x^{n+1}-\sum_{\color{red}{n=1}}^\infty na_{n}x^{n}+\sum_{\color{red}{n=1}}^\infty 2a_nx^n=-2a_0\\ \implies n(n+1)a_{n+1}-na_n+2a_n=-2a_0\\ \implies n(n+1)a_{n+1}=(n-2)a_n-2a_0\\ \implies\Large\boxed{a_{n+1}=\frac{(n-1)a_n-2a_0}{n(n+1)}} \]

OpenStudy (anonymous):

how are you combining your series sums when the x terms dont have the same exponents

OpenStudy (anonymous):

if a_0=0 \[a_{n+1}=\frac{(n-1)a_n}{n(n+1)}\]

OpenStudy (anonymous):

the first term should be x^n it is a type I am typing in tex and its rediculous to check my syntax!

OpenStudy (anonymous):

But I know I am right here

OpenStudy (anonymous):

okay its cool.

OpenStudy (anonymous):

darn.. \[ a_{n+1}=\frac{(n-2)a_n}{n(n+1)}\qquad\forall \;n\ne0 \] so, we have, for \[ n=1\quad a_2={-a_1\over2}\\ n=2\quad a_3=0\\ n=3\quad a_4=0\\\implies a_k=0\quad\forall\quad k\ge3\\ \Large y(x)=a_1x-{a_1\over 2}x^2=a_1x\left(1-{x\over2}\right)\\ y'=a_1-a_1x\\y''=-a_1 \]

OpenStudy (anonymous):

Oooooooooh okay I see where I went wrong now! I got it! thank you so much!!!

OpenStudy (anonymous):

yw phew

OpenStudy (anonymous):

still made a mistake.. here it is by hand..